### v1

# What Factors Affect LLMs and RLLMs in Financial Question Answering?

Abstract
Recently, the development of large language models (LLMs) and reasoning large language models (RLLMs) have gained considerable attention from many researchers. RLLMs enhance the reasoning capabilities of LLMs through Long Chain-of-Thought (Long CoT) processes, significantly improving the performance of LLMs in addressing complex problems. However, there are few works that systematically explore what methods can fully unlock the performance of LLMs and RLLMs within the financial domain. To investigate the impact of various methods on LLMs and RLLMs, we utilize five LLMs and three RLLMs to assess the effects of prompting methods, agentic frameworks, and multilingual alignment methods on financial question-answering tasks. Our research findings indicate: (1) Current prompting methods and agent frameworks enhance the performance of LLMs in financial question answering by simulating Long CoT; (2) RLLMs possess inherent Long CoT capabilities, which limits the effectiveness of conventional methods in further enhancing their performance; (3) Current advanced multilingual alignment methods primarily improve the multilingual performance of LLMs by extending the reasoning length, which yields minimal benefits for RLLMs. We hope that this study can serve as an important reference for LLMs and RLLMs in the field of financial question answering.
What Factors Affect LLMs and RLLMs in Financial Question Answering?
[…]
1 Introduction
Recently, large language models (LLMs) have significantly advanced the field of natural language processing (NLP), and more and more researchers utilize LLMs to solve complex tasks in various domains Zhou et al. (2023); Nie et al. (2024); Qin et al. (2024); Chang et al. (2024); Wang et al. (2025a). Furthermore, some researchers propose prompting-based and agentic frameworks to enhance the capabilities of LLMs Wei et al. (2022); Yao et al. (2023a); Zhang et al. (2024b); Yao et al. (2023b); Li et al. (2023); Hong et al. (2024), which further expands the capabilities of LLMs in tasks in different fields. Particularly, to enhance the reasoning capabilities of LLMs in addressing complex problems, researchers introduce reasoning large language models (RLLMs). The application of the Long Chain-of-Thought (Long CoT) fully unlocks the understanding ability of RLLMs Chen et al. (2025). Long CoT entails deeper reasoning, reflective analysis, and a more extensive exploration of logical structures Gandhi et al. (2025), which enables RLLMs to perform on par with experts in certain domains.
Researchers have explored the application of LLMs in financial question answering. Wu et al. (2023) present BloombergGPT, a 50 billion parameter LLM for financial domain. Fatemi and Hu (2024) propose a multi-agent framework to enhance the performance of LLMs in financial question answering. Xie et al. (2024) introduce a novel financial benchmark, FinBen, including 42 datasets spanning 24 financial tasks. Srivastava et al. (2024) evaluate LLMs in four financial tabular question-answering datasets. Xue et al. (2024) build a benchmark for financial multilingual multi-modal question answering named FAMMA, which is a challenging benchmark for multi-modal LLMs. Wang et al. (2025b) propose FinSage, which solves the financial filings question answering task through retrieval augmented generation (RAG).
Although LLMs have made significant progress in the field of financial question answering, existing work still mainly focuses on enhancing their performance within this domain. With the advancement of LLMs and RLLMs, the factors influencing their financial question-answering capabilities remain under-explored. Inspired by this gap, we systematically investigate the following issues: ($i$) the impact of prompting methods on LLMs/RLLMs; ($ii$) the influence of agentic frameworks on LLMs/RLLMs; and ($iii$) whether multilingual alignment methods can enhance the multilingual financial question-answering abilities of LLMs/RLLMs. We utilize five LLMs and three RLLMs, conducting detailed experiments across seven representative methods. Our code will be available at https://github.com/WPENGxs/LLM_RLLM_financial_analysis.
The mainly novel insights as follows:
  • Current prompting methods and agent frameworks enhance the performance of LLMs in financial question answering by simulating Long CoT. Effective prompting methods and agent frameworks primarily simulate Long CoT, enhancing LLM performance through extended reasoning lengths. This parallels the performance gains achieved by RLLMs through Long CoT, demonstrating that Long CoT represents a significant bottleneck for the current performance improvements of LLMs.
  • RLLMs possess inherent Long CoT capabilities, which limits the effectiveness of conventional methods in further enhancing their performance. Since RLLMs possess Long CoT capabilities, the conventional methods that are effective for LLMs cannot further enhance RLLM performance. We speculate that the key to improving RLLMs in the future lies in the introduction of more complex agent mechanisms. This would enable RLLMs to achieve excellent performance through deep reasoning while also standardizing previous outputs via the agent framework mechanisms.
  • Current advanced multilingual alignment methods primarily improve the multilingual performance of LLMs by extending the reasoning length, which yields minimal benefits for RLLMs. Current multilingual alignment methods depend on the inherent multilingual capabilities of the model, primarily by extending reasoning lengths to create mechanisms similar to Long CoT. This approach mirrors the gains observed with prompting methods and agentic frameworks in LLMs, thus it is not effective for RLLMs. Consequently, it becomes challenging for RLLMs to achieve further performance enhancements in multilingual contexts.
2 Background
Financial question answering involves inputting a question along with relevant context, resulting in the model generating the answer to the question. Specifically, for the multiple-choice question, the input comprises the question $\mathcal{Q}$, the available options $\mathcal{O}=\left\{l_{1},l_{2},...,l_{n}\right\}$, and the context $\mathcal{C}$, while the output consists of the selected options $\mathcal{A},\mathcal{A}\in\mathcal{O}$. For the open-ended question, the input includes the question $\mathcal{Q}$ and the context $\mathcal{C}$, with the output being the generated answer $\mathcal{A}$. These two types can be defined as follows:
$\mathcal{A}_{choice}=\underset{choice}{\operatorname{Model}}\left(\mathcal{Q},\mathcal{O},\mathcal{C}\right),$ (1)
$\mathcal{A}_{open}=\underset{open}{\operatorname{Model}}\left(\mathcal{Q},\mathcal{C}\right),$ (2)
where $\underset{choice}{\operatorname{Model}}\left(\cdot\right)$ presents the model for the multiple-choice question, $\underset{open}{\operatorname{Model}}\left(\cdot\right)$ presents the model for the open-ended question.
3 Exploration
To thoroughly investigate the impact of various methods on the performance of LLMs and RLLMs in financial question answering, this study examines three perspectives: prompting methods (§ 3.1), agentic frameworks (§ 3.2), and multilingual alignment methods (§ 3.3).
3.1 Exploration of prompting methods
Current research extensively demonstrates that different prompting methods can significantly alter the performance of LLMs Wei et al. (2022); Wang et al. (2023); Zhang et al. (2024b). The prompting method modifies certain aspects of the thinking prompt while ensuring the question remains intact, enabling LLMs to solve the question by following the structured steps of the thinking prompt.
We select three representative prompting methods to investigate the impact of varying prompt word changes on the performance of LLMs and RLLMs: Direct, Zero-shot CoT Kojima et al. (2022), and Plan-and-Solve Wang et al. (2023). These prompting methods can be defined as:
  • Direct: The direct prompting method enables LLMs to answer questions by inputting them directly, without the inclusion of the additional thinking prompt. This approach allows for an intuitive assessment of LLM performance and minimizes interference from redundant prompts.
  • Zero-shot CoT: The zero-shot CoT prompting employs the prompt “let’s think step by step” to stimulate the reasoning abilities of LLMs, enabling them to produce longer reasoning processes and, consequently, enhancing their performance.
  • Plan-and-Solve: The Plan-and-Solve prompting introduces a novel prompt “Let’s first understand the problem and devise a plan to solve the problem. Then, let’s carry out the plan to solve the problem step by step.” This addresses computational errors and enhances the quality of generated reasoning steps through a two-part decomposition.
3.2 Exploration of agentic frameworks
To further enhance the performance of LLMs, researchers have started to incorporate agentic systems. Specifically, by enabling LLMs to assume different roles and engage in continuous communication and cooperation, their performance can be significantly improved Luo et al. (2025).
We select two agentic frameworks, and aim to explore the impact of LLMs and RLLMs in different agent collaborations: Self-Refine Madaan et al. (2023) and S3 Agent Wang et al. (2024). These agentic frameworks can be defined as:
  • Self-Refine: Self-Refine provides feedback on the output using the same LLMs, thereby iteratively modifying the response and ultimately optimizing the answer to enhance performance. To save cost, we only use one iteration.
  • S3 Agent: S3 Agent proposes three agents of different perspectives: superficial expression, semantic information, and sentiment expression, to unlock the power of LLMs.
3.3 Exploration of multilingual alignment methods
Due to the varying proportions of different languages in the training corpus, the performance of LLMs typically differs across languages. Improving LLM performance in underrepresented languages has consistently been a key area of research. Current multilingual alignment methods primarily enhance the multilingual capabilities of LLMs through explicit prompts or translation.
We evaluate the performance of four standard multilingual approaches, including state-of-the-art cross-lingual prompting methods, on multilingual alignment: Direct, Translate-en Shi et al. (2023), and Cross-lingual Prompting Qin et al. (2023).
  • Direct: Direct prompting utilizes the English prompt and the local language question, which directly assess the multilingual capabilities.
  • Translate-en: Translate-en employs translation to convert the local language question into English, thereby preventing performance degradation of LLMs due to insufficient multilingual capabilities.
  • Cross-lingual Prompting (CLP): CLP enhances the reasoning abilities of LLMs on multilingual questions through a two-stage multilingual alignment based on prompts: (1) cross-lingual alignment prompting and (2) task-specific solver prompting.
Method Arithmetic Non-Arithmetic Overall Overall Easy Medium Hard Overall Easy Medium Hard Llama-3.1-8B-Instruct Grattafiori et al. (2024) Direct Grattafiori et al. (2024) 7.53 9.17 5.65 6.88 14.36 14.83 14.73 13.96 10.59 Zero-shot CoT Kojima et al. (2022) 8.18 11.35 4.24 7.18 22.18 23.30 20.00 22.52 14.44 Plan-and-Solve Wang et al. (2023) 11.16 13.53 9.18 9.58 22.06 25.84 17.89 21.84 16.04 Self-Refine Madaan et al. (2023) 13.67 18.55 8.12 11.67 25.28 26.27 23.68 25.45 18.86 S3 Agent Wang et al. (2024) 14.97 17.90 9.54 15.56 21.95 28.38 18.94 19.81 18.09 GPT-4o-mini Hurst et al. (2024) Direct Hurst et al. (2024) 22.32 30.56 18.02 14.67 30.80 37.71 29.47 27.70 26.11 Zero-shot CoT Kojima et al. (2022) 24.09 29.91 21.20 18.56 33.67 44.91 30.00 29.27 28.38 Plan-and-Solve Wang et al. (2023) 24.93 31.87 21.55 18.26 33.67 42.79 29.47 30.63 28.84 Self-Refine Madaan et al. (2023) 22.23 28.38 18.72 16.76 32.18 41.10 27.89 29.27 26.68 S3 Agent Wang et al. (2024) 21.86 28.82 18.37 15.26 31.49 41.10 28.42 27.70 26.16 Gemini-1.5-flash Team et al. (2024) Direct Team et al. (2024) 25.02 33.18 19.08 18.86 32.29 42.79 26.84 29.05 28.27 Zero-shot CoT Kojima et al. (2022) 24.74 32.53 19.08 18.86 32.52 42.79 27.89 29.05 28.22 Plan-and-Solve Wang et al. (2023) 25.76 33.18 22.61 18.26 33.33 42.37 28.94 30.40 29.15 Self-Refine Madaan et al. (2023) 21.86 28.16 16.60 17.66 29.65 36.44 25.78 27.70 25.34 S3 Agent Wang et al. (2024) 25.95 33.40 20.14 20.65 32.52 41.52 30.00 28.82 28.89 Qwen-2.5-32B Yang et al. (2024) Direct Yang et al. (2024) 27.90 35.80 22.61 21.55 35.17 43.64 31.57 32.20 31.15 Zero-shot CoT Kojima et al. (2022) 27.62 34.49 23.67 21.55 35.28 41.52 31.05 33.78 31.05 Plan-and-Solve Wang et al. (2023) 24.83 30.34 23.32 18.56 34.94 44.49 31.05 31.53 29.35 Self-Refine Madaan et al. (2023) 27.44 33.84 24.38 21.25 36.43 41.94 30.52 36.03 31.46 S3 Agent Wang et al. (2024) 29.20 36.68 24.38 23.05 35.28 43.64 34.73 31.08 31.92 DeepSeek-V3 Liu et al. (2024) Direct Liu et al. (2024) 35.06 42.13 29.32 30.23 43.21 51.27 35.78 42.11 38.71 Zero-shot CoT Kojima et al. (2022) 36.09 43.23 31.44 30.23 43.67 51.69 35.78 42.79 39.48 Plan-and-Solve Wang et al. (2023) 35.53 43.01 29.32 30.53 44.94 51.27 37.36 44.81 39.74 Self-Refine Madaan et al. (2023) 33.58 41.04 27.91 28.14 41.49 47.88 34.21 41.21 37.12 S3 Agent Wang et al. (2024) 35.34 41.92 30.38 30.53 42.52 51.27 35.78 40.76 38.56 DeepSeek-R1-Distill-Qwen-32B Guo et al. (2025) Direct Guo et al. (2025) 33.67 42.79 27.20 26.64 38.73 47.03 31.57 37.38 35.93 Zero-shot CoT Kojima et al. (2022) 34.69 42.79 30.03 27.54 38.27 49.15 30.52 35.81 36.29 Plan-and-Solve Wang et al. (2023) 33.86 42.79 26.85 27.54 36.20 45.76 30.52 33.55 34.91 Self-Refine Madaan et al. (2023) 28.37 35.15 24.02 22.75 33.79 38.13 27.36 34.23 30.79 S3 Agent Wang et al. (2024) 34.41 42.13 29.68 27.84 39.65 50.42 32.10 37.16 36.76 Qwen-3-14B Yang et al. (2025) Direct Yang et al. (2025) 33.30 43.23 28.97 23.35 38.62 49.57 32.63 35.36 35.68 Zero-shot CoT Kojima et al. (2022) 34.23 43.66 28.62 26.04 38.50 47.88 32.63 36.03 36.14 Plan-and-Solve Wang et al. (2023) 33.58 44.10 28.97 23.05 37.12 49.15 30.52 33.55 35.16 Self-Refine Madaan et al. (2023) 33.48 43.23 28.97 23.95 40.57 50.42 35.26 37.61 36.65 S3 Agent Wang et al. (2024) 34.04 43.44 30.03 24.55 40.68 51.69 33.68 37.83 37.01 O4-mini OpenAI (2025) Direct OpenAI (2025) 39.25 47.16 34.62 32.33 49.65 54.66 39.47 51.35 43.90 Zero-shot CoT Kojima et al. (2022) 39.62 46.50 33.21 35.62 50.57 54.23 42.63 52.01 44.52 Plan-and-Solve Wang et al. (2023) 39.72 46.50 34.27 35.02 49.54 53.81 40.52 51.12 44.11 Self-Refine Madaan et al. (2023) 38.97 46.06 31.44 35.62 49.08 52.11 39.47 51.57 43.49 S3 Agent Wang et al. (2024) 39.81 46.72 33.21 35.92 50.11 54.66 38.42 52.70 44.42
Table 1: The results of prompting and agentic methods. Bold number presents the best result for these methods on the current model. Underline number presents the second-best result for these methods on the current model. Light yellow presents the current model is LLM, Light green presents the current model is RLLM.
4 Experiments
4.1 Experimental Settings
We use the standard financial question answering benchmark FAMMA Xue et al. (2024) for experiments. FAMMA is a benchmark for financial multilingual multi-modal question answering, which includes English, Chinese, and French. Since current RLLMs are unimodal and cannot process multimodal information, we utilize the Basic Txt dataset in FAMMA, which converts multimodal data into textual information using OCR. The final dataset used for evaluation comprises 1,945 entries.
We conduct experiments on seven backbones, including five LLMs and three RLLMs: Meta-Llama-3.1-8B-Instruct Grattafiori et al. (2024), GPT-4o-mini Hurst et al. (2024), Gemini-1.5-flash Team et al. (2024), Qwen-2.5-32B Yang et al. (2024), DeepSeek-V3 Liu et al. (2024), DeepSeek-R1-Distill-Qwen-32B Guo et al. (2025), thing mode of Qwen-3-14B Yang et al. (2025), O4-mini OpenAI (2025). All LLMs and RLLMs use default settings.
4.2 Analysis of Prompting Methods
An effective prompting method significantly enhances the performance of LLMs. As shown in Table 1, advanced prompting method Plan-and-Solve outperforms other prompting and agentic methods on most LLMs. This suggests that employing a well-designed general prompting method effectively enhances the performance of LLMs, even if these methods are not originally proposed for financial question answering.
Larger LLMs exhibit increased robustness to various prompting methods. It is evident that for larger LLMs, such as Gemini-1.5-flash and DeepSeek-V3, the performance gap between different prompting methods is not significant, with performance fluctuations remaining under 1%. This indicates that LLMs with superior performance are less affected by input prompts, demonstrating greater robustness to variations in input.
Method Chinese French Overall (Chinese+French) Meta-Llama-3.1-8B-Instruct Grattafiori et al. (2024) Direct Grattafiori et al. (2024) 11.43 5.12 9.70 Translate-en Shi et al. (2023) 12.40 (+0.97) 6.41 (+1.29) 10.75 (+1.05) CLP Qin et al. (2023) 15.57 (+4.14) 5.12 (+0.00) 12.69 (+2.99) GPT-4o-mini Hurst et al. (2024) Direct Hurst et al. (2024) 24.81 13.46 21.69 Translate-en Shi et al. (2023) 25.79 (+0.98) 14.10 (+0.64) 22.57 (+0.88) CLP Qin et al. (2023) 27.98 (+3.17) 12.17 (-1.29) 23.63 (+1.94) Gemini-1.5-flash Team et al. (2024) Direct Team et al. (2024) 27.25 10.25 22.57 Translate-en Shi et al. (2023) 27.00 (-0.25) 9.61 (-0.64) 22.22 (-0.35) CLP Qin et al. (2023) 27.73 (+0.48) 9.61 (-0.64) 22.75 (+0.18) Qwen-2.5-32B Yang et al. (2024) Direct Yang et al. (2024) 30.17 13.46 25.57 Translate-en Shi et al. (2023) 32.36 (+2.19) 13.46 (+0.00) 27.16 (+1.59) CLP Qin et al. (2023) 31.14 (+0.97) 14.10 (+0.64) 26.45 (+0.88) DeepSeek-V3 Liu et al. (2024) Direct Liu et al. (2024) 38.44 22.43 34.03 Translate-en Shi et al. (2023) 38.44 (+0.00) 23.07 (+0.64) 34.21 (+0.18) CLP Qin et al. (2023) 40.87 (+2.43) 25.64 (+3.21) 36.68 (+2.65) DeepSeek-R1-Distill-Qwen-32B Guo et al. (2025) Direct Guo et al. (2025) 34.06 23.71 31.21 Translate-en Shi et al. (2023) 35.52 (+1.46) 21.15 (-2.56) 31.56 (+0.35) CLP Qin et al. (2023) 35.52 (+1.46) 19.23 (-4.48) 31.04 (-0.17) Qwen-3-14B Yang et al. (2025) Direct Yang et al. (2025) 37.22 19.87 32.45 Translate-en Shi et al. (2023) 38.19 (+0.97) 21.15 (+1.28) 33.50 (+1.05) CLP Qin et al. (2023) 35.76 (-1.46) 23.07 (+3.20) 32.27 (-0.18) O4-mini OpenAI (2025) Direct OpenAI (2025) 42.82 26.92 38.44 Translate-en Shi et al. (2023) 44.28 (+1.46) 26.28 (-0.64) 39.32 (+0.88) CLP Qin et al. (2023) 43.76 (+0.94) 26.92 (+0.00) 39.11 (+0.67)
Table 2: The results of multilingual alignment methods. Bold number presents the best result for these methods on the current model. Light yellow presents the current model is LLM, Light green presents the current model is RLLM. The performance of gains/drops relative to the Direct are highlight with green/red in the Table.
For RLLMs, competitive results can be achieved without the use of prompting methods. Benefiting from the Long CoT process, RLLMs can generate complex reasoning independently, without deliberate guidance from prompting methods. This renders prompting, which is effective for LLMs, less effective for RLLMs and may even diminish their performance. As shown in Table 1, Plan-and-Solve demonstrates poor performance on the three RLLMs.
4.3 Analysis of Agentic Frameworks
Smaller LLMs can derive greater benefits from the agentic framework. Smaller LLMs, such as Llama-3.1-8B-Instruct, can significantly improve their performance through the complex agentic framework, which is shown in Table 1. We speculate that this improvement may stem from the relatively weaker ability of smaller LLMs to understand and follow instructions. The standardized agentic framework helps mitigate the probability of the hallucination, thereby further enhancing overall performance.
A well-designed agentic framework is essential for further enhancing the performance of LLMs. For larger LLMs, directly employing the agentic framework does not lead to significant performance improvements. This may be due to the fact that these agentic frameworks are not specifically designed for financial question answering, and merely transferring them does not fully enhance LLM performance. To enhance performance in financial question answering, a well-designed agentic framework tailored to this task is necessary.
The performance gains of RLLMs within agentic frameworks primarily stem from effective agent collaboration. In RLLMs, agentic frameworks exhibit an improvement compared with prompting methods, particularly in Self-Refine and S3 Agent. This indicates that frameworks successful for LLMs can also be adapted for RLLMs. However, since RLLMs are developed through Long CoT, the performance enhancements associated with agentic frameworks, when employing longer thinking processes, are not pronounced. The benefits of these frameworks are more evident in their inherent design features, such as the reflective capabilities provided by Self-Refine and the multi-faceted thinking fostered by S3 Agent.
Figure 1: The performance of Qwen-2.5-32B and DeepSeek-R1-Distill-Qwen-32B.
4.4 Analysis of Multilingual Alignment Methods
The Translate-en method can enhance multilingual performance for LLMs to some extent; however, the improvement is not substantial. As indicated in Table 2, the Translate-en method yields gains across most models compared to the Direct approach, but these gains remain insignificant. This may be attributed to the fact that Translate-en aligns multilingual tasks to a single language. Consequently, the overall reasoning length for LLMs does not increase, resulting in a lack of depth in thinking, which in turn limits the significant enhancement of performance.
Extending the reasoning length significantly enhances the performance of LLMs. For most LLMs, employing the CLP method yields better performance than the Translate-en approach. This indicates that CLP not only aligns multiple languages to the target language but also further improves performance by strengthening the Long CoT of LLMs. Additionally, it is noteworthy that in the CLP method, the performance gain for Chinese is more pronounced than that for French. Since the initial stage alignment performance of CLP is also influenced by the inherent multilingual capabilities of the LLMs, insufficient multilingual ability may limit the performance gains achievable through CLP.
RLLMs demonstrate self-alignment capabilities for multilingual questions. In RLLMs, the effectiveness of various alignment methods, such as Translate-en and CLP, is limited. It is posited that RLLMs can achieve results comparable to those of CLP through Long CoT. This suggests that the introduction of CLP may induce overthinking in RLLMs, resulting in a decline in performance. While translation does yield some performance improvement, the enhancement is not significant.
4.5 The benefits of Long CoT cannot be offset by prompting methods and agentic frameworks
We compare the performance of Qwen-2.5-32B and DeepSeek-R1-Distill-Qwen-32B, where DeepSeek-R1-Distill-Qwen-32B is a Qwen model distilled from DeepSeek-R1, as illustrated in Figure 1. It is evident that DeepSeek-R1-Distill-Qwen-32B outperforms the base model across various methods, demonstrating an average improvement of 3.95%. This indicates that the gains from Long CoT significantly enhance the reasoning capabilities of the base model after RLLMs distillation. Consequently, it is challenging for the base model to surpass RLLMs equipped with Long CoT capabilities through advanced frameworks. While it is possible to narrow this gap using a more complex framework, we argue that this approach may lead to token consumption of LLMs approaching that of RLLMs with the same base model.
Figure 2: Statistics of average output token consumption for questions of different difficulty in LLMs/RLLMs, only counting the correct output. The line chart shows the performance of LLMs/RLLMs in the Direct method.
Model Direct Zero-shot CoT Plan-and-Solve Llama-3.1-8B-Instruct 181.50 378.34 478.02 gpt-4o-mini 198.87 395.58 493.43 gemini-1.5-flash 205.36 221.17 234.45 Qwen-2.5-32B 283.00 369.96 364.24 DeepSeek-V3 364.13 395.04 420.60 DeepSeek-R1-Distill-Qwen-32B 1,721.73 1750.14 1858.72 Qwen-3-14B 1,696.88 1701.82 1665.87
Table 3: Statistics of average correct output token consumption in current LLMs/RLLMs.
Figure 3: Performance of different scale and thinking mode in Qwen-3. The line chart in (a) shows the performance of Qwen-3 in the Direct method. The histogram in (a) shows the average correct output tokens at different scales. The histogram in (b) shows the performance of Qwen-3 in different methods.
4.6 Longer reasoning processes generally lead to better performance
To explore the relationship between output length and performance in LLMs and RLLMs, we conduct a statistical analysis of output tokens. The statistics presented in Table 3 indicate that within the same model, more robust methods require a greater number of tokens. This suggests a positive correlation between performance and token usage in the current LLMs, effective prompting techniques enhance performance by simulating a Long CoT. Statistic of Direct in Figure 2 reveals that, in LLMs, longer reasoning processes with improved performance. This suggests that sufficient cognitive processing benefits LLMs and their current limitations primarily stem from a constrained thinking process.
In contrast, for RLLMs, the output of Long CoT indicates that the length of the CoT is not the primary determinant of performance. For instance, DeepSeek-R1-Distill-Qwen-32B generates an excessive number of tokens for easy questions, yet this does not lead to significant performance enhancements compared with Qwen-3-14B. In easy questions, DeepSeek-R1-Distill-Qwen-32B achieves 44.23%, Qwen-3-14B achieves 45.38% in Acc. This phenomenon may arise from the overthinking in RLLMs, which fails to yield performance gains while substantially increasing computational costs. Moving forward, dynamically adjusting CoT length according to the complexity of the input problem will represent a critical area of research.
4.7 Performance follows LLM parameter and output increase
We further analyze the performance and token consumption of Qwen-3 across various scales and the thinking mode, with the final results illustrated in Figure 3. The data reveal that the performance of LLMs adheres to the scaling law Kaplan et al. (2020), larger models exhibit enhanced performance and longer output lengths. Additionally, activating the thinking mode provides a significant boost, yielding an average improvement of 35% for the Qwen-3-14B model.
Figure 4: Results of CLP on questions of different difficulty. The transparent color indicates the performance of the Direct method.
4.8 CLP significantly enhances the performance of LLMs and RLLMs when addressing complex problems
To investigate how the CLP method improves the multilingual capabilities of various LLMs and RLLMs, we analyze model performance across different difficulty levels, as illustrated in Figure 4. The results demonstrate that CLP effectively boosts model performance on challenging problems and substantially enhances multilingual proficiency. This alignment method proves to be beneficial even for RLLMs.
We posit that, for LLMs, CLP optimally enhances performance through the reasoning length and the alignment of diverse languages. For smaller LLMs, gains are observed across varying levels of difficulty. Conversely, for larger LLMs, CLP primarily improves performance on difficult tasks. While CLP can yield benefits for RLLMs through multilingual alignment, it may also lead to overthinking simpler problems, resulting in less noticeable overall gains.
Local Language English Chinese French Target Language Chinese French English French English Chinese Meta-Llama-3.1-8B-Instruct 13.78 8.05 15.57 9.97 5.12 5.76 GPT-4o-mini 29.24 27.93 27.98 25.30 12.17 12.17 Gemini-1.5-flash 30.40 30.69 27.73 28.46 9.61 12.17 Qwen-2.5-32B 32.00 31.78 31.14 31.63 14.10 14.74 DeepSeek-V3 42.38 41.79 40.87 40.38 25.64 23.71 DeepSeek-R1-Distill-Qwen-32B 35.41 36.57 35.52 35.27 19.23 21.15 Qwen-3-14B 38.60 38.24 35.76 38.92 23.07 20.51
Table 4: The results of CLP. Light green presents the performance of CLP is better than the Direct method, Light red presents the performance of CLP is worse than the Direct method.
4.9 CLP must select the most effective alignment language to maximize performance
We explore the performance differences across various target languages within the CLP method, with the final results presented in Table 4. The findings indicate that different target languages yield varying performance gains, and directly aligning local languages to English through CLP is often not the most effective strategy. Furthermore, different models utilize distinct optimal target languages identified by the CLP method, which may be attributed to varying language capabilities among the models. Although some researchers propose CLSP Qin et al. (2023) and AutoCAP Zhang et al. (2024a). CLSP investigates the integration of CLP and self-consistency methods to bridge the gap between different languages. AutoCAP automatically selects languages and automatic weight allocation to obtain the best language combination. This approach incurs substantial operational costs. Therefore, the selection of the best target language based on the multilingual capabilities of each model requires further exploration.
5 Related Work
With the development of LLMs, more and more researchers focus on the application of LLM in various fields. Improving the performance of LLMs in downstream tasks has consistently been a focal point of researchers’ efforts. Wei et al. (2022) explore the impact of Chain-of-Thought on LLMs and find that forcing thinking of LLMs can significantly improve performance. Inspired by this, researchers further explored the improvement of LLMs. Yao et al. (2023a) propose Tree-of-Thought (ToT), which enhances LLM by considering multiple reasoning paths and self-evaluating. Yin et al. (2023) introduce Exchange-of-Thought (EoT) to enable cross-model communication. In EoT, LLMs could enhance their performance through different network topologies. Zhang et al. (2024b) employs Multi-Perspective Verification and Wrong Information Utilization to prevent LLMs from repeating the same mistakes, thereby significantly enhancing their performance.
Some researchers have introduced complex agentic frameworks to enhance performance by standardizing the processes of LLMs. Yao et al. (2023b) present a novel agent framework ReAct. ReAct can track the interactions of LLMs with the environment and make decisions about the next steps. Li et al. (2023) propose a novel communicative agent framework that enables LLMs to understand tasks through role-playing. Hong et al. (2024) introduce Standardized Operating Procedures (SOPs) into frameworks and propose MetaGPT, which enhances the ability of software engineering in LLMs. Wang et al. (2024) propose S3 Agent to improve the performance of LLMs on multi-modal problems by introducing multiple perspectives information.
6 Conclusion
This paper is the first to explore the performance of various methods for LLMs and RLLMs in the financial question-answering task, focusing on why these methods are effective for improvement in LLMs. We conduct experiments on five LLMs and three RLLMs, revealing the critical role of Long CoT in the performance of language models. Our results indicate that the goal of these methods is to simulate Long CoT. We hope these findings deepen our understanding of the Long CoT and serve as a valuable reference in financial question-answering.
Limitations
Due to computational cost constraints, we primarily selected representative methods for our experiments. These methods may not fully reflect the performance of the majority of techniques in financial question answering. To prevent data leakage, we utilized the latest benchmark, FAMMA, for our experiments. However, we cannot guarantee that similar data exists for some questions in the internet data. The challenge of LLMs fitting data requires further exploration in future work.