### v4

# What Factors Affect LLMs and RLLMs in Financial Question Answering?

Abstract
Recently, large language models (LLMs) and reasoning large language models (RLLMs) have gained considerable attention from many researchers. RLLMs enhance the reasoning capabilities of LLMs through Long Chain-of-Thought (Long CoT) processes, significantly improving the performance of LLMs in addressing complex problems. However, there are few works that systematically explore what methods can fully unlock the performance of LLMs and RLLMs within the financial domain. To investigate the impact of various methods on LLMs and RLLMs, we utilize five LLMs and four RLLMs to assess the effects of prompting methods, agentic frameworks, and multilingual alignment methods on financial question-answering tasks. Our research findings indicate: (1) Current prompting methods and agent frameworks enhance the performance of LLMs in financial question answering by simulating Long CoT; (2) RLLMs possess inherent Long CoT capabilities, which limits the effectiveness of conventional methods in further enhancing their performance; (3) Current advanced multilingual alignment methods primarily improve the multilingual performance of LLMs by extending the reasoning length, which yields minimal benefits for RLLMs. Additionally, we discuss strategies for enhancing the performance of LLMs and RLLMs in financial question answering, which may serve as a inspiration for future improvements. We hope that this study can serve as an important reference for LLMs and RLLMs in the field of financial question answering.
What Factors Affect LLMs and RLLMs in Financial Question Answering?
[…]
1 Introduction
Recently, large language models (LLMs) have significantly advanced the field of natural language processing (NLP), and more and more researchers utilize LLMs to solve complex tasks in various domains (zhou2023survey; nie2024survey; qin2024large; chang2024survey; peng2025large). Furthermore, some researchers propose prompting-based and agentic frameworks to enhance the capabilities of LLMs (wei2022chain; yao2023tree; zhang-etal-2024-wrong; yao2023react; li2023camel; hong2024metagpt), which further expands the capabilities of LLMs in tasks in different fields. Particularly, to enhance the reasoning capabilities of LLMs in addressing complex problems, researchers introduce reasoning large language models (RLLMs). The application of the Long Chain-of-Thought (Long CoT) fully unlocks the understanding ability of RLLMs (chen2025towards). Long CoT entails deeper reasoning, reflective analysis, and a more extensive exploration of logical structures (gandhi2025cognitive), which enables RLLMs to perform on par with experts in certain domains.
Researchers have explored the application of LLMs in financial question answering. wu2023bloomberggpt present BloombergGPT, a 50 billion parameter LLM for financial domain. 10.1145/3677052.3698686 propose a multi-agent framework to enhance the performance of LLMs in financial question answering. xie2024finben introduce a novel financial benchmark, FinBen, including 42 datasets spanning 24 financial tasks. srivastava2024evaluating evaluate LLMs in four financial tabular question-answering datasets. xue2024famma build a benchmark for financial multilingual multi-modal question answering named FAMMA, which is a challenging benchmark for multi-modal LLMs. wang2025finsage propose FinSage, which solves the financial filings question answering task through retrieval augmented generation (RAG).
Although LLMs have made significant progress in the field of financial question answering, existing work still mainly focuses on enhancing their performance within this domain. With the advancement of LLMs and RLLMs, the factors influencing their financial question-answering capabilities remain under-explored. Inspired by this gap, we systematically investigate the following issues: ($i$) the impact of prompting methods on LLMs/RLLMs; ($ii$) the influence of agentic frameworks on LLMs/RLLMs; and (iiiiii) whether multilingual alignment methods can enhance the multilingual financial question-answering abilities of LLMs/RLLMs. We utilize five LLMs and four RLLMs, conducting detailed experiments across seven representative methods.
The mainly novel insights as follows:
  • Current prompting methods and agent frameworks enhance the performance of LLMs in financial question answering by simulating Long CoT. Effective prompting methods and agent frameworks primarily simulate Long CoT, enhancing LLM performance through extended reasoning lengths. This parallels the performance gains achieved by RLLMs through Long CoT, demonstrating that Long CoT represents a significant bottleneck for the current performance improvements of LLMs.
  • RLLMs possess inherent Long CoT capabilities, which limits the effectiveness of conventional methods in further enhancing their performance. Since RLLMs possess Long CoT capabilities, the conventional methods that are effective for LLMs cannot further enhance RLLM performance. We speculate that the key to improving RLLMs in the future lies in the introduction of more complex agent mechanisms. This would enable RLLMs to achieve excellent performance through deep reasoning while also standardizing previous outputs via the agent framework mechanisms.
  • Current advanced multilingual alignment methods primarily improve the multilingual performance of LLMs by extending the reasoning length, which yields minimal benefits for RLLMs. Current multilingual alignment methods depend on the inherent multilingual capabilities of the model, primarily by extending reasoning lengths to create mechanisms similar to Long CoT. This approach mirrors the gains observed with prompting methods and agentic frameworks in LLMs, thus it is not effective for RLLMs. Consequently, it becomes challenging for RLLMs to achieve further performance enhancements in multilingual contexts.
2 Background
Financial question answering involves inputting a question along with relevant context, resulting in the model generating the answer to the question. Specifically, for the multiple-choice question, the input comprises the question $\mathcal{Q}$, the available options $\mathcal{O}=\left\{l_{1},l_{2},...,l_{n}\right\}$, and the context $\mathcal{C}$, while the output consists of the selected options $\mathcal{A},\mathcal{A}\in\mathcal{O}$. For the open-ended question, the input includes the question $\mathcal{Q}$ and the context $\mathcal{C}$, with the output being the generated answer $\mathcal{A}$. These two types can be defined as follows:
$\mathcal{A}_{choice}=\underset{choice}{\operatorname{Model}}\left(\mathcal{Q},\mathcal{O},\mathcal{C}\right),$ (1)
$\mathcal{A}_{open}=\underset{open}{\operatorname{Model}}\left(\mathcal{Q},\mathcal{C}\right),$ (2)
where $\underset{choice}{\operatorname{Model}}\left(\cdot\right)$ presents the model for the multiple-choice question, $\underset{open}{\operatorname{Model}}\left(\cdot\right)$ presents the model for the open-ended question.
3 Exploration
To thoroughly investigate the impact of various methods on the performance of LLMs and RLLMs in financial question answering, this study examines three perspectives: prompting methods (§ 3.1), agentic frameworks (§ 3.2), and multilingual alignment methods (§ 3.3).
3.1 Exploration of prompting methods
Current research extensively demonstrates that different prompting methods can significantly alter the performance of LLMs (wei2022chain; wang2023plan; zhang-etal-2024-wrong). The prompting method modifies certain aspects of the thinking prompt while ensuring the question remains intact, enabling LLMs to solve the question by following the structured steps of the thinking prompt.
We select three representative prompting methods to investigate the impact of varying prompt word changes on the performance of LLMs and RLLMs: Direct, Zero-shot CoT (kojima2022large), and Plan-and-Solve (wang2023plan). These prompting methods can be defined as:
  • Direct: The direct prompting method enables LLMs to answer questions by inputting them directly, without the inclusion of the additional thinking prompt. This approach allows for an intuitive assessment of LLM performance and minimizes interference from redundant prompts.
  • Zero-shot CoT: The zero-shot CoT prompting employs the prompt “let’s think step by step” to stimulate the reasoning abilities of LLMs, enabling them to produce longer reasoning processes and, consequently, enhancing their performance.
  • Plan-and-Solve: The Plan-and-Solve prompting introduces a novel prompt “Let’s first understand the problem and devise a plan to solve the problem. Then, let’s carry out the plan to solve the problem step by step.” This addresses computational errors and enhances the quality of generated reasoning steps through a two-part decomposition.
3.2 Exploration of agentic frameworks
To further enhance the performance of LLMs, researchers have started to incorporate agentic systems. Specifically, by enabling LLMs to assume different roles and engage in continuous communication and cooperation, their performance can be significantly improved (luo2025large).
We select two agentic frameworks, and aim to explore the impact of LLMs and RLLMs in different agent collaborations: Self-Refine (madaan2023self) and S3 Agent (10.1145/3690642). These agentic frameworks can be defined as:
  • Self-Refine: Self-Refine provides feedback on the output using the same LLMs, thereby iteratively modifying the response and ultimately optimizing the answer to enhance performance. To save cost, we only use one iteration.
  • S3 Agent: S3 Agent proposes three agents of different perspectives: superficial expression, semantic information, and sentiment expression, to unlock the power of LLMs. We made minor modifications to the S3 Agent for financial question-answering tasks while preserving the three core components intact.
3.3 Exploration of multilingual alignment methods
Due to the varying proportions of different languages in the training corpus, the performance of LLMs typically differs across languages. Improving LLM performance in underrepresented languages has consistently been a key area of research. Current multilingual alignment methods primarily enhance the multilingual capabilities of LLMs through explicit prompts or translation.
We evaluate the performance of four standard multilingual approaches, including state-of-the-art cross-lingual prompting methods, on multilingual alignment: Direct, Translate-en (shi2023language), and Cross-lingual Prompting (qin-etal-2023-cross).
  • Direct: Direct prompting utilizes the English prompt and the local language question, which directly assess the multilingual capabilities.
  • Translate-en: Translate-en employs translation to convert the local language question and context into English, thereby preventing performance degradation of LLMs due to insufficient multilingual capabilities.
  • Cross-lingual Prompting (CLP): CLP enhances the reasoning abilities of LLMs on multilingual questions through a two-stage multilingual alignment based on prompts: (1) cross-lingual alignment prompting and (2) task-specific solver prompting.
Model Method Arithmetic Non-Arithmetic Overall Overall Easy Medium Hard Overall Easy Medium Hard Llama-3.1-8B-Instruct Direct grattafiori2024llama 10.88 12.66 9.89 9.28 23.44 22.45 32.10 20.27 16.50 (grattafiori2024llama) Zero-shot CoT kojima2022large 13.20 16.15 10.95 11.07 31.72 32.62 34.73 29.95 21.49 Plan-and-Solve wang2023plan 16.09 18.99 14.84 13.17 29.77 32.62 30.52 27.92 22.21 Self-Refine madaan2023self 17.95 22.70 13.07 15.56 32.06 34.74 35.26 29.27 24.26 S3 Agent 10.1145/3690642 19.62 22.27 15.54 19.46 30.80 37.28 32.63 26.57 24.62 GPT-4o-mini Direct hurst2024gpt 32.46 41.70 30.38 21.55 42.52 50.42 50.52 34.90 36.96 (hurst2024gpt) Zero-shot CoT kojima2022large 34.13 41.04 34.27 24.55 45.17 56.35 52.10 36.26 39.07 Plan-and-Solve wang2023plan 36.18 44.97 33.92 26.04 45.74 56.77 51.05 37.61 40.46 Self-Refine madaan2023self 33.30 41.26 33.21 22.45 42.06 52.11 45.78 35.13 37.22 S3 Agent 10.1145/3690642 30.88 38.42 28.26 22.75 41.72 51.69 46.84 34.23 35.73 Gemini-1.5-flash Direct team2024gemini 37.20 47.16 35.33 25.14 45.28 58.05 49.47 36.71 40.82 (team2024gemini) Zero-shot CoT kojima2022large 37.39 46.72 36.04 25.74 46.09 58.89 51.57 36.93 41.28 Plan-and-Solve wang2023plan 37.76 46.50 39.57 24.25 47.12 55.50 55.26 39.18 41.95 Self-Refine madaan2023self 34.41 42.35 33.92 23.95 42.52 52.96 50.00 33.78 38.04 S3 Agent 10.1145/3690642 37.20 45.85 34.62 27.54 44.59 57.20 50.52 35.36 40.51 Qwen-2.5-32B Direct yang2024qwen2 41.86 50.65 40.98 30.53 48.62 59.32 56.31 39.63 44.88 (yang2024qwen2) Zero-shot CoT kojima2022large 42.32 50.21 40.98 32.63 50.80 59.32 60.00 42.34 46.11 Plan-and-Solve wang2023plan 39.44 44.97 43.46 28.44 49.77 60.59 55.78 41.44 44.06 Self-Refine madaan2023self 41.58 49.56 42.75 29.64 49.65 57.62 54.73 43.24 45.19 S3 Agent 10.1145/3690642 42.51 51.52 41.34 31.13 48.85 61.44 56.31 38.96 45.34 DeepSeek-V3 Direct liu2024deepseek 56.27 65.50 58.65 41.61 62.06 73.72 69.47 52.70 58.86 (liu2024deepseek) Zero-shot CoT kojima2022large 56.18 63.97 59.71 42.51 61.83 71.18 70.52 53.15 58.71 Plan-and-Solve wang2023plan 55.16 63.97 56.18 42.21 63.33 70.76 71.05 56.08 58.81 Self-Refine madaan2023self 53.86 61.79 55.83 41.31 59.54 69.06 64.73 52.25 56.40 S3 Agent 10.1145/3690642 54.69 61.57 56.89 43.41 59.42 71.18 66.31 50.22 56.81 DeepSeek-R1-Distill-Qwen-32B Direct guo2025deepseek 50.41 60.48 50.17 36.82 57.12 65.25 64.73 49.54 53.41 (guo2025deepseek) Zero-shot CoT kojima2022large 51.25 60.91 51.94 37.42 56.55 68.64 63.15 47.29 53.62 Plan-and-Solve wang2023plan 51.16 60.91 50.53 38.32 55.86 66.52 62.63 47.29 53.26 Self-Refine madaan2023self 44.65 52.18 47.70 31.73 52.06 57.20 60.00 45.94 47.96 S3 Agent 10.1145/3690642 51.25 59.60 51.94 39.22 58.04 69.49 65.26 48.87 54.29 GPT-OSS-20B Direct agarwal2025gpt 50.04 61.57 48.40 35.62 57.93 60.59 67.89 52.25 53.57 (agarwal2025gpt) Zero-shot CoT kojima2022large 50.04 60.69 51.59 34.13 56.89 64.40 68.94 47.74 53.11 Plan-and-Solve wang2023plan 49.67 59.82 49.46 35.92 55.28 63.98 59.47 48.87 52.18 Self-Refine madaan2023self 48.27 56.55 48.76 36.52 57.58 69.49 61.57 49.54 52.44 S3 Agent 10.1145/3690642 52.00 61.79 51.23 39.22 58.62 68.22 63.68 51.35 54.96 Qwen-3-14B Direct yang2025qwen3 54.88 66.59 58.65 35.62 59.08 72.03 70.52 47.29 56.76 (yang2025qwen3) Zero-shot CoT kojima2022large 55.72 68.34 56.18 38.02 59.19 69.91 70.52 48.64 57.27 Plan-and-Solve wang2023plan 54.97 66.81 58.65 35.62 57.81 71.18 68.94 45.94 56.24 Self-Refine madaan2023self 54.41 65.93 56.89 36.52 58.27 71.61 66.84 47.52 56.14 S3 Agent 10.1145/3690642 54.13 65.06 57.24 36.52 58.50 71.18 67.36 47.97 56.09 O4-mini Direct o4-mini 61.30 70.08 65.01 46.10 70.22 77.54 75.78 63.96 65.29 (o4-mini) Zero-shot CoT kojima2022large 62.60 69.86 64.66 50.89 71.37 77.11 77.89 65.54 66.52 Plan-and-Solve wang2023plan 61.67 69.86 64.31 48.20 72.41 77.96 77.89 67.11 66.47 Self-Refine madaan2023self 61.20 69.21 60.77 50.59 70.34 75.00 74.73 65.99 65.29 S3 Agent 10.1145/3690642 61.76 69.65 62.19 50.59 71.49 76.69 75.78 66.89 66.11
Table 1: The results of prompting and agentic methods. Bold number presents the best result for these methods on the current model. Underline number presents the second-best result for these methods on the current model. Light yellow presents the current model is LLM, Light green presents the current model is RLLM.
4 Experiments
4.1 Experimental Settings
We use the standard financial question answering benchmark FAMMA (xue2024famma) for experiments. FAMMA is a benchmark for financial multilingual multi-modal question answering, which includes English, Chinese, and French. Since current RLLMs are unimodal and cannot process multimodal information, we utilize the Basic Txt dataset in FAMMA, which converts multimodal data into textual information using OCR. The final dataset used for evaluation comprises 1,945 entries.
We conduct experiments on nine backbones, including five LLMs and four RLLMs: Meta-Llama-3.1-8B-Instruct (grattafiori2024llama), GPT-4o-mini (hurst2024gpt), Gemini-1.5-flash (team2024gemini), Qwen-2.5-32B (yang2024qwen2), DeepSeek-V3 (liu2024deepseek), DeepSeek-R1-Distill-Qwen-32B (guo2025deepseek), GPT-OSS-20B (agarwal2025gpt), think mode of Qwen-3-14B (yang2025qwen3), O4-mini (o4-mini). All open questions are evaluated using GPT-4o-mini based on correct answers. Due to computational constraints, all models were evaluated using a single run.
4.2 Analysis of Prompting Methods
An effective prompting method significantly enhances the performance of LLMs. As shown in Table 1, advanced prompting method Plan-and-Solve outperforms other prompting and agentic methods on most LLMs. This suggests that employing a well-designed general prompting method effectively enhances the performance of LLMs, even if these methods are not originally proposed for financial question answering.
Larger LLMs exhibit increased robustness to various prompting methods. It is evident that for larger LLMs, such as Gemini-1.5-flash and DeepSeek-V3, the performance gap between different prompting methods is not significant, with performance fluctuations remaining under 2%. This indicates that LLMs with superior performance are less affected by input prompts, demonstrating greater robustness to variations in input.
Performance Direct Translate-en (shi2023language) Self-Refine (madaan2023self) S3 Agent (10.1145/3690642) CLP (qin-etal-2023-cross) Meta-Llama-3.1-8B-Instruct (grattafiori2024llama) Chinese 16.54 16.54 (+0.00) 18.97 (+2.43) 19.22 (+2.68) 23.35 (+6.81) French 15.38 16.66 (+1.28) 13.46 (-1.92) 21.15 (+5.77) 17.30 (+1.92) Overall (Chinese+French) 16.22 16.57 (+0.35) 17.46 (+1.24) 19.75 (+3.53) 21.69 (+5.47) GPT-4o-mini (hurst2024gpt) Chinese 33.09 34.54 (+1.45) 34.79 (+1.70) 31.63 (-1.46) 38.19 (+5.10) French 28.84 30.76 (+1.92) 28.84 (+0.00) 32.69 (+3.85) 32.05 (+3.21) Overall (Chinese+French) 31.92 33.50 (+1.58) 33.15 (+1.23) 31.92 (+0.00) 36.50 (+4.58) Gemini-1.5-flash (team2024gemini) Chinese 40.38 38.68 (-1.70) 36.73 (-3.65) 39.65 (-0.73) 37.22 (-3.16) French 27.56 29.48 (+1.92) 26.92 (-0.64) 27.56 (+0.00) 29.48 (+1.92) Overall (Chinese+French) 36.86 36.15 (-0.71) 34.03 (-2.83) 36.33 (-0.53) 35.09 (-1.77) Qwen-2.5-32B (yang2024qwen2) Chinese 41.84 44.52 (+2.68) 44.28 (+2.44) 42.82 (+0.98) 43.30 (+1.46) French 29.48 29.48 (+0.00) 33.97 (+4.49) 35.89 (+6.41) 35.25 (+5.77) Overall (Chinese+French) 38.44 40.38 (+1.94) 41.44 (+3.00) 40.91 (+2.47) 41.09 (+2.65) DeepSeek-V3 (liu2024deepseek) Chinese 57.42 58.63 (+1.21) 55.96 (-1.46) 57.66 (+0.24) 61.31 (+3.89) French 50.00 50.00 (+0.00) 47.43 (-2.57) 44.23 (-5.77) 57.69 (+7.69) Overall (Chinese+French) 55.37 56.26 (+0.89) 53.61 (-1.76) 53.96 (-1.41) 60.31 (+4.94) DeepSeek-R1-Distill-Qwen-32B (guo2025deepseek) Chinese 48.41 49.87 (+1.46) 46.47 (-1.94) 52.06 (+3.25) 50.85 (+2.44) French 49.35 42.94 (-6.41) 44.87 (-1.16) 49.35 (+0.00) 52.56 (+3.21) Overall (Chinese+French) 48.67 47.97 (-0.70) 46.03 (-2.64) 51.32 (+2.65) 51.32 (+2.65) GPT-OSS-20B (agarwal2025gpt) Chinese 50.36 51.33 (+0.97) 50.60 (+0.24) 54.50 (+4.14) 54.76 (+4.40) French 51.28 46.15 (-5.13) 46.15 (-5.13) 55.76 (+4.48) 45.51 (-5.77) Overall (Chinese+French) 50.61 49.91 (-0.70) 49.38 (-1.23) 54.85 (+4.24) 52.21 (+1.60) Qwen-3-14B (yang2025qwen3) Chinese 57.66 56.69 (-0.97) 55.23 (-2.43) 55.47 (-2.19) 53.77 (-3.89) French 48.71 50.00 (+1.29) 42.3 (-6.41) 48.71 (+0.00) 51.28 (+2.57) Overall (Chinese+French) 55.20 54.85 (-0.35) 51.67 (-3.53) 53.61 (-1.59) 53.08 (-2.12) O4-mini (o4-mini) Chinese 65.45 65.93 (+0.48) 64.23 (-1.22) 67.63 (+2.18) 64.54 (-0.91) French 62.17 61.53 (-0.64) 60.25 (-1.92) 61.53 (-0.64) 63.46 (+1.29) Overall (Chinese+French) 64.55 64.72 (+0.17) 63.13 (-1.42) 65.69 (+1.14) 64.24 (-0.31)
Table 2: The results of multilingual alignment methods. Bold number presents the best result for these methods on the current model. Light yellow presents the current model is LLM, Light green presents the current model is RLLM. The performance of gains/drops relative to the Direct are highlight with green/red in the Table.
For RLLMs, competitive results can be achieved without the use of prompting methods. Benefiting from the Long CoT process, RLLMs can generate complex reasoning independently, without deliberate guidance from prompting methods. This renders prompting, which is effective for LLMs, less effective for RLLMs and may even diminish their performance. As shown in Table 1, Plan-and-Solve demonstrates poor performance on the four RLLMs.
4.3 Analysis of Agentic Frameworks
Smaller LLMs can derive greater benefits from the agentic framework. Smaller LLMs, such as Llama-3.1-8B-Instruct, can significantly improve their performance through the complex agentic framework, which is shown in Table 1. We speculate that this improvement may stem from the relatively weaker ability of smaller LLMs to understand and follow instructions. The standardized agentic framework helps mitigate the probability of the hallucination, thereby further enhancing overall performance.
A well-designed agentic framework is essential for further enhancing the performance of LLMs. For larger LLMs, directly employing the agentic framework does not lead to significant performance improvements. This may be due to the fact that these agentic frameworks are not specifically designed for financial question answering, and merely transferring them does not fully enhance LLM performance. To enhance performance in financial question answering, a well-designed agentic framework tailored to this task is necessary.
The performance gains of RLLMs within agentic frameworks primarily stem from effective agent collaboration. In RLLMs, agentic frameworks exhibit an improvement compared with prompting methods, particularly in S3 Agent for DeepSeek-R1-Distill-Qwen-32B. This indicates that frameworks successful for LLMs can also be adapted for RLLMs. However, since RLLMs are developed through Long CoT, the performance enhancements associated with agentic frameworks, when employing longer thinking processes, are not pronounced. The benefits of these frameworks are more evident in their inherent design features, such as the reflective capabilities provided by Self-Refine and the multi-faceted thinking fostered by S3 Agent.
4.4 Analysis of Multilingual Alignment Methods
The Translate-en method can enhance multilingual performance for LLMs to some extent; however, the improvement is not substantial. As indicated in Table 2, the Translate-en method yields gains across most models compared to the Direct approach, but these gains remain insignificant. This may be attributed to the fact that Translate-en aligns multilingual tasks to a single language. Consequently, the overall reasoning length for LLMs does not increase, resulting in a lack of depth in thinking, which in turn limits the significant enhancement of performance.
Figure 1: The performance of Qwen-2.5-32B and DeepSeek-R1-Distill-Qwen-32B.
Extending the reasoning length significantly enhances the performance of LLMs. For most LLMs, employing the CLP method yields better performance than the Translate-en approach. Additionally, the agentic framework (Self-Refine and S3 Agent) enhances the performance of LLMs to some degree. This indicates that CLP not only aligns multiple languages to the target language but also further improves performance by strengthening the Long CoT of LLMs. Since the initial stage alignment performance of CLP is also influenced by the inherent multilingual capabilities of the LLMs, insufficient multilingual ability may limit the performance gains achievable through CLP.
RLLMs demonstrate self-alignment capabilities for multilingual questions. In RLLMs, the effectiveness of various alignment methods, such as Translate-en and CLP, is limited. It is posited that RLLMs can achieve results comparable to those of CLP through Long CoT. This suggests that the introduction of CLP may induce overthinking in RLLMs, resulting in a decline in performance. While translation does yield some performance decline, the reduction is not significant.
4.5 The benefits of Long CoT cannot be offset by prompting methods and agentic frameworks
We compare the performance of Qwen-2.5-32B and DeepSeek-R1-Distill-Qwen-32B, where DeepSeek-R1-Distill-Qwen-32B is a Qwen-2.5 model distilled from DeepSeek-R1, as illustrated in Figure 1. It is evident that DeepSeek-R1-Distill-Qwen-32B outperforms the base model across various methods, demonstrating an average improvement of 7.4%. This indicates that the gains from Long CoT significantly enhance the reasoning capabilities of the base model after RLLMs distillation. Consequently, it is challenging for the base model to surpass RLLMs equipped with Long CoT capabilities through advanced frameworks. While it is possible to narrow this gap using a more complex framework, we argue that this approach may lead to token consumption of LLMs approaching that of RLLMs with the same base model.
Figure 2: Statistics of average output token consumption for questions of different difficulty in LLMs/RLLMs. The line chart shows the performance of LLMs/RLLMs in the Direct method.
Model Direct Zero-shot CoT Plan-and-Solve Meta-Llama-3.1-8B-Instruct 464.36 973.27 1131.86 gpt-4o-mini 249.77 430.41 531.83 gemini-1.5-flash 286.32 296.03 313.42 Qwen-2.5-32B 368.91 470.99 476.74 DeepSeek-V3 425.65 450.58 493.32 DeepSeek-R1-Distill-Qwen-32B 1,972.60 2006.39 2196.63 Qwen-3-14B 2,087.49 2120.97 2211.74
Table 3: Statistics of average output token consumption in current LLMs/RLLMs.
4.6 Longer reasoning processes generally lead to better performance
To explore the relationship between output length and performance in LLMs and RLLMs, we conduct a statistical analysis of output tokens. The statistics presented in Table 3 indicate that within the same model, more robust methods require a greater number of tokens. This suggests a positive correlation between performance and token usage in the current LLMs, effective prompting techniques enhance performance by simulating a Long CoT. Statistic of Direct in Figure 2 reveals that, in most LLMs, longer reasoning processes with improved performance. This suggests that sufficient cognitive processing benefits LLMs and their current limitations primarily stem from a constrained thinking process.
Figure 3: Performance of different scale and thinking mode in Qwen-3. The line chart in (a) shows Direct performance of Qwen-3, while the histogram in (a) shows average output tokens across scales. The histogram in (b) shows performance in different methods.
In contrast, for RLLMs, the output of Long CoT indicates that the length of the CoT is not the primary determinant of performance. For instance, DeepSeek-R1-Distill-Qwen-32B generates an excessive number of tokens for easy questions, yet this does not lead to significant performance enhancements compared with Qwen-3-14B. In easy questions, DeepSeek-R1-Distill-Qwen-32B achieves 53.41%, Qwen-3-14B achieves 56.76% in Acc. This phenomenon may arise from the overthinking in RLLMs (chen2024not), which fails to yield performance gains while substantially increasing computational costs. Moving forward, dynamically adjusting CoT length according to the complexity of the input problem will represent a critical area of research.
4.7 Performance improves as the parameters of LLMs increase
We further analyze the performance and token consumption of Qwen-3 across various scales and the thinking mode, with the final results illustrated in Figure 3. The data reveal that the performance of LLMs adheres to the scaling law (kaplan2020scaling), larger models exhibit enhanced performance and longer output lengths. This suggests that more powerful LLMs engage in more extensive reasoning processes, resulting in superior thinking chain capabilities.
Additionally, we examine the performance of the Qwen-3-14B model with thinking mode activated and deactivated. Activating the thinking mode yields a significant improvement, with an average increase of 16.9%. This indicates that RLLMs can achieve superior performance through extended reasoning processes.
Figure 4: Error statistics for Qwen-3-14B under different thinking modes, considering cases where thinking mode is correct but non-thinking mode fails.
Source Language English Chinese French Target Language Chinese French English French English Chinese Meta-Llama-3.1-8B-Instruct 19.88 14.00 23.35 13.38 17.30 18.58 GPT-4o-mini 39.47 38.53 38.19 35.03 32.05 28.84 - Gemini-1.5-flash 41.14 41.50 37.22 38.92 29.48 30.76 Qwen-2.5-32B 45.57 44.99 43.30 42.82 35.25 33.97 DeepSeek-V3 62.55 60.88 61.31 59.36 57.69 53.84 DeepSeek-R1-Distill-Qwen-32B 50.79 51.45 50.85 48.90 52.56 51.28 Qwen-3-14B 57.91 57.03 53.77 58.39 51.28 50.00
Table 4: The results of CLP. presents the performance of CLP is better than the Direct method, presents the performance of CLP is worse than the Direct method.
4.8 Understanding the Advantages of Long CoT
To elucidate the advantages conferred by Long CoT, we utilize GPT-4o-mini to systematically analyze the error types observed in Qwen-3-14B under both thinking and non-thinking modes, as illustrated in Figure 4.
The analysis reveals that, in the absence of the thinking mode, the reasoning error constitutes the predominant source of performance degradation. In contrast, activation of the thinking mode substantially mitigates such errors, highlighting the efficacy of Long CoT in enhancing the reasoning capabilities of RLLMs. Furthermore, RLLMs demonstrate superior comprehension of problem statements compared to LLMs, as evidenced by the fact that approximately one-third of LLM errors are attributable to misinterpretation of questions and underlying assumptions. These findings collectively underscore the critical role of Long CoT in advancing the overall performance of financial question answering models.
4.9 It’s important to select the optimal alignment language in CLP
We explore the performance differences across various target languages within the CLP method, with the final results presented in Table 4. The findings indicate that different target languages yield varying performance gains, and directly aligning local languages to English through CLP is often not the most effective strategy. Utilizing various language alignments has demonstrated differing impacts on performance (wang-etal-2025-x), likely attributable to LLMs’ comprehension of distinct language families.
Furthermore, different models utilize distinct optimal target languages identified by the CLP method, which may be attributed to varying language capabilities among the models. Although some researchers propose CLSP (qin-etal-2023-cross) and AutoCAP (zhang2024autocap). These approaches incur substantial operational costs. Therefore, the selection of the best target language based on the multilingual capabilities of each model requires further exploration.
5 Related Work
With the development of LLMs, more and more researchers focus on the application of LLM in various fields. Improving the performance of LLMs in downstream tasks has consistently been a focal point of researchers’ efforts. wei2022chain explore the impact of Chain-of-Thought on LLMs and find that forcing thinking of LLMs can significantly improve performance. Inspired by this, researchers further explored the improvement of LLMs. yao2023tree propose Tree-of-Thought (ToT), which enhances LLM by considering multiple reasoning paths and self-evaluating. yin-etal-2023-exchange introduce Exchange-of-Thought (EoT) to enable cross-model communication. In EoT, LLMs could enhance their performance through different network topologies. zhang-etal-2024-wrong employs Multi-Perspective Verification and Wrong Information Utilization to prevent LLMs from repeating the same mistakes, thereby significantly enhancing their performance.
Some researchers have introduced complex agentic frameworks to enhance performance by standardizing the processes of LLMs. yao2023react present a novel agent framework ReAct. ReAct can track the interactions of LLMs with the environment and make decisions about the next steps. li2023camel propose a novel communicative agent framework that enables LLMs to understand tasks through role-playing. hong2024metagpt introduce Standardized Operating Procedures (SOPs) into frameworks and propose MetaGPT, which enhances the ability of software engineering in LLMs. 10.1145/3690642 propose S3 Agent to improve the performance of LLMs on multi-modal problems by introducing multiple perspectives information.
6 Conclusion
In this work, we systematically investigate the factors influencing the performance of LLMs and RLLMs in financial question answering. Through extensive experiments on multiple models and methods, we find that the effectiveness of prompting strategies, agentic frameworks, and multilingual alignment approaches for LLMs largely stems from their ability to simulate longer reasoning chains (Long CoT). RLLMs, equipped with inherent Long CoT capabilities, show limited improvement from these conventional methods, highlighting a performance bottleneck. Our analysis suggests that future advances for RLLMs may require more sophisticated agentic mechanisms and dynamic reasoning processes. We hope these findings deepen our understanding of the Long CoT and serve as a valuable reference in financial question-answering.
Limitations
Due to computational cost constraints, we primarily selected representative methods for our experiments. These methods may not fully reflect the performance of the majority of techniques in financial question answering. To prevent data leakage, we utilized the latest benchmark, FAMMA, for our experiments. However, we cannot guarantee that similar data exists for some questions in the internet data. The challenge of LLMs fitting data requires further exploration in future work.
Acknowledgments
This work was supported by Fund for the Development of Science and Technology (FDCT) of Macau under Grant 0095/2023/RIA2.
Appendix A The Use of Large Language Models (LLMs)
We declare that only LLMs were utilized to polish the English of this paper.
Appendix B Discussion
While considerable research has focused on enhancing the performance of LLMs across various domains, our experiments indicate that the key to improving LLM performance may lie in strengthening their long-term thinking capabilities, which aligns with the overarching objective of RLLMs. Consequently, we find that currently effective prompting methods and agentic frameworks for LLMs do not yield significant improvements for RLLMs. It is important to note that the agentic framework employed in our study was not specifically tailored for financial question answering. Thus, agentic frameworks that integrate RLLMs with techniques such as retrieval-augmented generation (RAG) and are well-designed for financial question answering may present a promising solution.
Furthermore, the Long CoT approach constrains the potential for enhancing performance of RLLMs through agentic frameworks. The advent of agentic reinforcement learning (Agentic RL) enables LLMs and RLLMs to function not only as standalone foundation models but also to be incorporated into complex reasoning and decision-making processes (luo2025agent; zhang2025landscape). Developing a suitable agentic workflow and further training LLMs and RLLMs through agentic reinforcement learning represents an intriguing avenue for improving model performance.
Model Overall (non-thinking) Overall (thinking) Qwen-3-4B 20.35 40.56 Qwen-3-8B 27.66 45.70 Qwen-3-14B 38.09 56.76
Table 5: The performance of the Direct method by Qwen-3-4B/8B/14B in non-thinking and thinking mode.
Appendix C Performance scaling in Qwen-3 non-thinking/thinking mode
We evaluate the performance of Qwen-3 in Direct method under non-thinking/thinking mode, and the final comparison table is shown in Table 5. It can be seen that even small-scale models can benefit from Long CoT.
Method Arithmetic Non-Arithmetic Overall Llama-3.1-8B-Instruct Direct 2.22 12.06 7.76 Zero-shot CoT 6.66 6.89 6.79 Plan-and-Solve 6.66 6.89 6.79 Self-Refine 8.88 6.89 7.76 S3 Agent 2.22 8.62 5.82 GPT-4o-mini Direct 11.11 10.34 10.67 Zero-shot CoT 13.33 17.24 15.53 Plan-and-Solve 17.77 12.06 14.56 Self-Refine 20.00 15.51 17.47 S3 Agent 15.55 3.44 8.73 DeepSeek-R1-Distill-Qwen-32B Direct 26.66 32.75 30.09 Zero-shot CoT 20.00 37.93 30.09 Plan-and-Solve 20.00 32.75 27.18 Self-Refine 17.77 32.75 26.21 S3 Agent 24.44 41.37 33.98 Qwen-3-14B Direct 22.22 37.93 31.06 Zero-shot CoT 33.33 41.37 37.86 Plan-and-Solve 24.44 46.55 36.89 Self-Refine 24.44 36.20 31.06 S3 Agent 22.22 34.48 29.12
Table 6: The results of prompting and agentic methods in FAMMA-LivePro. Bold number presents the best result for these methods on the current model. Light yellow presents the current model is LLM, Light green presents the current model is RLLM.
Appendix D The additional experiment in FAMMA-LivePro
We also conduct a small-scale experiment on FAMMA-LivePro, an additional benchmark independent of the main experiment in this paper, which consists of 103 original questions written by senior financial practitioners. The final experimental results are shown in Table 6. The final results are broadly similar to the conclusions in our paper, so we can draw the same conclusions as the original paper.
Direct Zero-shot CoT Plan-and-Solve Meta-Llama-3.1-8B-Instruct Cost (OpenRouter, Output \$0.05/M) 0.045 0.095 0.110 Overall Acc. 16.50 21.49 22.21 GPT-4o-mini Cost (OpenAI, output \$0.60/M) 0.291 0.502 0.621 Overall Acc. 16.96 39.07 40.46 Qwen-2.5-32B Cost (DeepInfra, output \$0.39/M) 0.280 0.357 0.362 Overall Acc. 44.88 46.11 44.06 DeepSeek-V3 Cost (DeepInfra, output \$0.89/M) 0.737 0.780 0.854 Overall Acc. 58.56 58.71 58.81 DeepSeek-R1-Distill-Qwen-32B Cost (OpenRouter, output \$0.29/M) 1.113 1.132 1.239 Overall Acc. 53.41 53.62 53.26 Qwen-3-14B Cost (OpenRouter, output \$0.24/M) 0.974 0.990 1.032 Overall Acc. 56.76 57.27 56.24
Table 7: The budget results of different methods. Light yellow presents the current model is LLM, Light green presents the current model is RLLM.
Appendix E RLLMs exhibit cost stability across prompting strategies
We analyze the cost of some tokens, and the final results are shown in Table 7. The price of the closed-source model follows the official price, while the price of the open-source model is based on OpenRouter and DeepInfra.
We observe a striking divergence in the cost dynamics between LLMs and RLLMs when subjected to different prompting methods. Specifically, the inference cost of LLMs exhibits high volatility depending on the applied method, whereas RLLMs maintain cost stability. For instance, increasing for Meta-Llama-3.1-8B-Instruct and GPT-4o-mini alongside corresponding accuracy improvements. Conversely, RLLMs display minimal cost variations across prompting methods.